Caskey Inc. is experiencing a period of growth. Dividends are expected to grow at a rate of 13.00% for the next two years and 5.00% thereafter. Yesterday the corporation paid a dividend of $1.18. If the required rate of return is 13.00%, what is the intrinsic value of the stock?
Answer #1

D1=(1.18*1.13)=1.3334

D2=(1.3334*1.13)=1.506742

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(1.506742*1.05)/(0.13-0.05)

=19.7759887

Hence intrinsic value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.3334/1.13+1.506742/1.13^2+19.7759887/1.13^2

=$17.8475